If x= 2 and y = 4 ,then x ^y/x + y^x/yFind x and y

Grade:9

1 Answers

Sheetal kumari
6 years ago
x^y/x+y^x/y= we know the value of x=2 and y=4 2^4/2+4^2/4= 4+2=6 hope you are finding the value of thus equation and not the value of x and y as it is given please kindly approved me